OpenStudy (mikeymaximum):

Which fraction is between 3/6 and 1? A. 1/3 B. 5/12 C.1/11 D.9/10

OpenStudy (mikeymaximum):

@mizanin

OpenStudy (a_clan):

Find decimal values for each.

OpenStudy (anonymous):

is anybody good at Spanish? or do you know somebody

OpenStudy (mikeymaximum):

Well i'm kinda good at it

OpenStudy (anonymous):

i really need help if you dont mind

OpenStudy (mikeymaximum):

okay just put it down right here

OpenStudy (anonymous):

ok sure

OpenStudy (mikeymaximum):

next time you should go to english

OpenStudy (mikeymaximum):

@mizanin can u help please

OpenStudy (mikeymaximum):

@mizanin can you help plaese

OpenStudy (mikeymaximum):

@Crissy15

OpenStudy (mikeymaximum):

@pooja195

pooja195 (pooja195):

ill tell u the same thing as @A.clan find the decimal value for each

OpenStudy (mikeymaximum):

Hibthis is not really mikeymaximum this acatally @bigi but my brother let me use his profile my profile is not woking so he said to use his now you remember @bigi right

OpenStudy (mikeymaximum):

I don't get what your saying

OpenStudy (mikeymaximum):

@pooja195 i don't get what your saying

OpenStudy (crissy15):

convert each fraction into a decimal 1/3= 5/12= 1/11= 9/10= start then from there call me

OpenStudy (mikeymaximum):

THIS IS THE REAL @MikeyMaximum AND I SAY THAT THE ANSWER IS A FRACTION THAT BIGGER THAN 1/2